# Ore extensions of multiplier Hopf coquasigroups

### v1

@@ 3 Ore extensions of multiplier Hopf coquasigroups
Next, we will present the main result of this section, which gives the necessary and sufficient conditions for the Ore extension of a multiplier Hopf coquasigroup to be a multiplier Hopf coquasigroup.
Theoerm 3.4. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. Assume that $\tau$ is a surjective endomorphism on $A$. Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is a MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
@@
$\tau(a)=\chi(a_{(1)})a_{(2)};$ (3.2)
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)};$ (3.3)
@@
(2) The $\tau-derivation$ $\delta$ satisfies the relation
$\Delta(\delta(a))=\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})$ (3.4)
@@
Proof. The proof will be conducted in three steps.
step1. Comultiplication.
Similar to the contruction in [14], assume that the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau.\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$. Then the algebra homomorphism $\Delta$ preserves the relation $ya=\tau(a)y+\delta(a)$ for any $a\in A$, i.e.,
@@
$=\tau({a_{(1)}})y\otimes{a_{(2)}}b+\delta({a_{(1)}})\otimes{a_{(2)}}b+r{a_{(1)}}\otimes\tau({a_{(2)}})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
@@
$=\big{(}\tau({a_{(1)}})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
$\quad+\big{(}r{a_{(1)}}{r^{-1}}\otimes\delta({a_{(2)}})\big{)}(r\otimes b)+\big{(}r{a_{(1)}}{r^{-1}}\otimes\tau({a_{(2)}})\big{)}\big{(}r\otimes\delta(b)\big{)}+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
and
@@
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}$
$\quad+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
So we have that $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following formulas hold:
@@
By the definition of $\delta\otimes\iota$ and $\iota\otimes\delta$ in [14, Lemma 2.5] and the extensions of $\tau\otimes\iota$, the above conditions change to the following:
$\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
@@
[…]
@@
for any $a,b\in A$. Because $\tau$ is nondegenerate, for $b$ there are $a_{i}$ , $b_{i}$ such that $\sum{{a_{i}}}\tau\left({{b_{i}}}\right)=b$. Hence, the above definition is changed to the following:
$\chi(a)b$ $=m(id\otimes S)\Big{(}\sum\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\Big{(}\sum(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)\Big{(}(\tau\otimes id)(b_{i}\otimes 1)\big{(}(\tau\otimes id)\Delta(a)\big{)}\Big{)}\bigg{)}$
$=m(id\otimes S)\bigg{(}\sum(a_{i}\otimes 1)(\tau\otimes id)\Big{(}(b_{i}\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}\bigg{)}.$
Hence $\chi(a)b\in A$ for any $a,b\in A$. In addition, for any $c\in A$,
$\chi(a)(bc)$ $=m(id\otimes S)\big{(}(bc\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=m(id\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
$=b\big{(}m(id\otimes S)(c\otimes 1)(\tau\otimes id)\Delta(a)\big{)}$
@@
[…]
$\chi(a)b=m(id\otimes S)\Big{(}(b\otimes 1)(\tau\otimes id)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
and since $b$ is arbitrary, this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}.$ Therefore, $\chi(A)\subseteq k$. We can regard $\chi$ as a mapping $\chi:A\rightarrow k$. Considering that $\tau$ is an algebra endomorphism and
@@
[…]
$(id\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
[…]
@@
By applying the formula $\chi\left(a\right)=\varepsilon\left(\tau\left(a\right)\right)$, $\tau$ can be derived from $\chi$,
@@
$=\varepsilon\big{(}\tau(a_{(1)})\big{)}S(a_{(2)(1)})a_{(2)(2)}$
$=\varepsilon(1)\tau(a)=\tau(a).$
[…]
@@
[…]
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}=\tau(a_{(1)})\otimes a_{(2)},$
$\Delta\big{(}\tau(a)\big{)}$ $=\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$=Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}=Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
[…]
Therefore, we can obtain that the comultiplication of $A$ (i.e., the comultiplication on $M(A)$ ) can be extended to an algebra homomorphism $\Delta:M(A)[y;\tau,\delta]\rightarrow M(A)[y;\tau,\delta]\otimes M(A)[y;\tau,\delta]$, and $\Delta$ does not necessarily have to be coassociative.
@@
Assume that $A[y;\tau,\delta]$ already has the comultiplication defined as in the above step. In what follows, we will prove the existence of a counit denoted by $\overline{\varepsilon}$ extending $\varepsilon$ from $A$(in fact $M(A)$) to the Ore extension $M(A)[y;\tau,\delta]$. Define $\overline{\varepsilon}(y)=0$. We need to verify the equation
$\overline{\varepsilon}(y)\varepsilon(x)=\varepsilon\big{(}\tau(x)\big{)}\overline{\varepsilon}(y)+\varepsilon\big{(}\delta(x)\big{)}$
for any $x\in M(A)$, i.e., $\varepsilon\big{(}\delta(x)\big{)}=0$. In fact, we only have to prove the above formula is true for any $a\in A$ in the sense that
$0=\varepsilon\big{(}\delta(xa)\big{)}=\varepsilon\big{(}\delta(x)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}.$
[…]
$\delta(a)$ $=m(id\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
$=m(id\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
@@
Therefore, $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0$
This proves $\varepsilon\big{(}\delta(a)\big{)}=0$ , for any $a\in A$. Furthermore, since
$0=\varepsilon\big{(}\delta(xa)\big{)}$ $=\varepsilon\big{(}\delta(x)a+\tau(x)\delta(a)\big{)}$
$=\varepsilon\big{(}\delta(a)\big{)}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\varepsilon\big{(}\delta(x)\big{)}$
$=\varepsilon\big{(}\delta(x)\big{)}\Big{(}\varepsilon(a)+\varepsilon\big{(}\tau(x)\big{)}\Big{)},$
it follows that for any $x\in M(A)$, $\varepsilon\big{(}\delta(x)\big{)}=0$ holds true. Consequently, $\varepsilon$ can be extended to $M(A)[y;\tau,\delta]$.
For all $\sum a_{i}y^{i},\sum b_{j}y^{j}\in A[y;\tau,\delta]$, it is easy to prove
$(\overline{\varepsilon}\otimes\iota)\big{(}\Delta(\sum a_{i}y^{i})(1\otimes\sum b_{j}y^{j})\big{)}$ $=(\sum a_{i}y^{i})(\sum b_{j}y^{j}),$
$(\iota\otimes\overline{\varepsilon})\big{(}(\sum b_{j}y^{j}\otimes 1)\Delta(\sum a_{i}y^{i})\big{)}$ $=(\sum b_{j}y^{j})(\sum a_{i}y^{i}).$
@@
[…]
@@
$S(a)S(y)=S(y)S\big{(}\alpha(a)\big{)}+S\big{(}\delta(a)\big{)}$ (3.10)
[…]
$1\otimes\delta(a)$ $=(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
@@
Applying the formula mentioned above, we can obtain
@@
$=(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau({a_{(1)}})yb\otimes{a_{(2)}}+\tau({a_{(1)}})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
@@
$=S\big{(}\tau({a_{(1)}})yb\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}y\otimes{a_{(2)(2)}}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
$\quad+S(r{a_{(1)}}b)\otimes\delta({a_{(2)(1)}})\otimes{a_{(2)(2)}}+S(r{a_{(1)}}b)\otimes r{a_{(2)(1)}}\otimes\delta({a_{(2)(2)}})\Big{)}$
@@
$\quad+S(b)S(r)S\big{(}\tau({a_{(1)}})\big{)}{a_{(2)(1)}}r\otimes{a_{(2)(2)}}y+S\big{(}\delta(a)b\big{)}{a_{(2)(1)}}\otimes{a_{(2)(2)}}$
@@
=(3.11)${=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$ )
@@
[…]
@@
$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)},$
for any $a\in M(A).$
[…]

### v2

@@ 3 Ore extensions of multiplier Hopf coquasigroups
We now proceed to state the principal theorem of this section, providing necessary and sufficient conditions under which the Ore extension of a multiplier Hopf coquasigroup again admits a multiplier Hopf coquasigroup structure.
Theorem 3.4. Let $A$ and $A[y;\tau,\delta]$ be regular multiplier Hopf coquasigroups. Assume that $\tau$ is a surjective nondegenerate algebra endomorphism on $A$. Then the regular multiplier Hopf coquasigroup $A[y;\tau,\delta]$ is an MHC-Ore extension of $A$ if and only if there is a group-like element $r\in M(A)$ such that the following conditions hold:
@@
$\tau(a)=\chi(a_{(1)})a_{(2)}=(\chi\otimes\iota)\Delta(a)$ (3.2)
@@
$(2)$ The following equations hold
$\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}=Ad_{r}(a_{(1)})\chi(a_{(2)(1)})\otimes a_{(2)(2)}=\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}$ (3.3)
@@
(3)(3)The $\tau$-derivation $\delta$ satisfies the relation
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}$ (3.4)
for any $a\in A$.
Proof. The proof proceeds in two parts, we begin with the necessity.
1. Proof of necessity.
Analogous to the construction in [16], the comultiplication $\Delta$ can be extended from $A$ to $A[y;\tau,\delta]$ by $\Delta(y)=y\otimes 1+r\otimes y$. Then the algebra homomorphism $\Delta$ preserves the relation $ya=\tau(a)y+\delta(a)$ for any $a\in A$, i.e.,
@@
$\overset{\bigstar}{=}\tau(a_{(1)})y\otimes a_{(2)}b+\delta(a_{(1)})\otimes a_{(2)}b+ra_{(1)}\otimes\tau(a_{(2)})\tau(b)y+r{a_{(1)}}\otimes\delta({a_{(2)}})b$
@@
$=\big{(}\tau(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b)(y\otimes 1)+\big{(}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)})\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
$\quad+\big{(}Ad_{r}(a_{(1)})\otimes\delta(a_{(2)})\big{)}(r\otimes b)+\big{(}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)})\big{)}\big{(}r\otimes\delta(b)\big{)}$
$\quad+\big{(}\delta(a_{(1)})\otimes a_{(2)}\big{)}(1\otimes b),$
where the identity $ra_{(1)}\otimes\tau(a_{(2)}b)y=ra_{(1)}\otimes\tau(a_{(2)})\tau(b)y$ at step $\bigstar$ follows from the fact that $\tau$ is a nondegenerate algebra homomorphism. Furthermore,
@@
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)y\big{)}$
$\quad+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b)$
$=\Delta\big{(}\tau(a)\big{)}(1\otimes b)(y\otimes 1)+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\tau(b)\big{)}(1\otimes y)$
$\quad+\Delta\big{(}\tau(a)\big{)}\big{(}r\otimes\delta(b)\big{)}+\Delta\big{(}\delta(a)\big{)}(1\otimes b).$
Therefore, the equality $\Delta(y)\Delta(a)=\Delta(\tau(a))\Delta(y)+\Delta(\delta(a))$ holds if and only if the following three conditions are satisfied: for all $b\in A$,
@@
Using the properties of $\delta\otimes\iota$ and $\iota\otimes\delta$ from [16, Lemma 2.5], along with the extensions of $\tau\otimes\iota$, we reformulate the above conditions as follows:
$\Delta\big{(}\tau(a)\big{)}=(\tau\otimes\iota)\Delta(a)=\tau(a_{(1)})\otimes a_{(2)},$ (3.5)
@@
$\Delta\big{(}\delta(a)\big{)}=(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}.$
The final condition is specified by equation $(\ref{3.4})$.
We now demonstrate that equations $(\ref{3.5})$ and $(\ref{3.6})$ imply the existence of a character $\chi$ such that both $(\ref{3.2})$ and $(\ref{3.3})$ are valid. Following the approach in the proof of Proposition 3.4 in [16], we define a map $\chi:A\rightarrow R(A)$ by
@@
for any $a,b\in A$. Since $\tau$ is nondegenerate, for any $b$ there exist elements $a_{i}$ $b_{i}$ such that $\sum_{i=1}^{n}{{a_{i}}}\tau\left({{b_{i}}}\right)=b$. Hence, the expression above can be rewritten as:
$\chi(a)b$ $=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}\big{(}a_{i}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}$
$=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}(a_{i}\otimes 1)\big{(}\tau(b_{i})\otimes 1\big{)}\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}$
$=m(\iota\otimes S)\bigg{(}\sum_{i=1}^{n}(a_{i}\otimes 1)\Big{(}(\tau\otimes\iota)(b_{i}\otimes 1)\big{(}(\tau\otimes\iota)\Delta(a)\big{)}\Big{)}\bigg{)}$
$=m(\iota\otimes S)\Big{(}\sum_{i=1}^{n}(a_{i}\otimes 1)(\tau\otimes\iota)\big{(}(b_{i}\otimes 1)\Delta(a)\big{)}\Big{)}.$
It follows that $\chi(a)b\in A$ for all $a,b\in A$. In addition, for all $c\in A$,
$\chi(a)(bc)$ $=m(\iota\otimes S)\big{(}(bc\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
$=m(\iota\otimes S)\big{(}(b\otimes 1)(c\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
$=b\big{(}m(\iota\otimes S)(c\otimes 1)(\tau\otimes\iota)\Delta(a)\big{)}$
@@
This shows that $\chi(a)$ is a right multiplier on $A$ for any $a\in A$. According to the formula $(\ref{3.5})$, the properties of $S$ imply that for any $a,b\in A$
$\chi(a)b\overset{(\ref{3.5})}{=}m(\iota\otimes S)\Big{(}(b\otimes 1)\Delta\big{(}\tau(a)\big{)}\Big{)}=b\varepsilon\big{(}\tau(a)\big{)},$
this proves that $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}$. Therefore, $\chi(A)\subseteq k$ and $\chi$ can be regarded as a map $\chi:A\rightarrow k$. Considering that $\tau$ is an algebra endomorphism and
@@
which implies that $\chi$ is a character on $A$ and is nondegenerate.
Now, for $\tau(a)\in A$, equation $(\ref{3.5})$ gives
$(\iota\otimes\Delta)\Delta\big{(}\tau(a)\big{)}=\tau(a_{(1)})\otimes a_{(2)(1)}\otimes a_{(2)(2)}.$
Then by $(\ref{2.1})$ we have
@@
Using the formula $\chi(a)=\varepsilon\big{(}\tau(a)\big{)}$, we can express $\tau$ in terms of $\chi$,
@@
$=\varepsilon\big{(}\tau(a_{(1)})S(a_{(2)(1)})\big{)}a_{(2)(2)}$
$\overset{(\ref{3.7})}{=}\varepsilon(1)\tau(a)=\tau(a).$
This establishes equation $(\ref{3.2})$. By substituting $\tau(a)=\chi(a_{(1)})a_{(2)}$(as established above) into both sides of equations $(\ref{3.5})$ and $(\ref{3.6})$, it ensues that
@@
These two equations are exactly $(\ref{3.3})$.
2. Proof of sufficiency.
The sufficiency part of the proof is carried out in three steps.
step 1. Comultiplication.
Under conditions $(\ref{3.2})$-$(\ref{3.4})$, we proceed to show that $(\ref{3.5})$ and $(\ref{3.6})$ hold. We have
$\Delta\big{(}\tau(a)\big{)}$ $\overset{(\ref{3.2})}{=}\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$\overset{(\ref{3.3})}{=}\chi(a_{(1)(1)})a_{(1)(2)}\otimes a_{(2)}\overset{(\ref{3.2})}{=}\tau(a_{(1)})\otimes a_{(2)},$
$\Delta\big{(}\tau(a)\big{)}$ $\overset{(\ref{3.2})}{=}\Delta\big{(}\chi(a_{(1)})a_{(2)}\big{)}=\chi(a_{(1)})a_{(2)(1)}\otimes a_{(2)(2)}$
$\overset{(\ref{3.3})}{=}Ad_{r}(a_{(1)})\otimes\chi(a_{(2)(1)})a_{(2)(2)}\overset{(\ref{3.2})}{=}Ad_{r}(a_{(1)})\otimes\tau(a_{(2)}).$
These equations confirm that $(\ref{3.5})$ and $(\ref{3.6})$ hold. Together with the necessity part of the proof, we conclude that $\Delta(y)\Delta(a)=\Delta\big{(}\tau(a)\big{)}\Delta(y)+\Delta\big{(}\delta(a)\big{)}$ is satisfied. The comultiplication $\Delta:A[y;\tau,\delta]\rightarrow M(A[y;\tau,\delta]\otimes A[y;\tau,\delta])$ is defined as above, note that it is not necessarily coassociative.
@@
Assume that $A[y;\tau,\delta]$ is equipped with the comultiplication defined in Step 1. Suppose further that the counit $\varepsilon$ can be extended to $A[y;\tau,\delta]$, with $\varepsilon(y)=0$, and remains an algebra homomorphism. Then the relation $ya=\tau(a)y+\delta(a)$ must be preserved under $\varepsilon$, i.e.,
$\varepsilon(y)\varepsilon(a)=\varepsilon\big{(}\tau(a)\big{)}\varepsilon(y)+\varepsilon\big{(}\delta(a)\big{)}$
for all $a\in A$. Substituting $\varepsilon(y)=0$, we find that such an extension exists if and only if
$\varepsilon\big{(}\delta(a)\big{)}=0.$
We now show that $\varepsilon\big{(}\delta(a)\big{)}=0$ indeed holds for all $a\in A$. Applying the map $m(\iota\otimes\varepsilon)$ to $(\ref{3.4})$, we obtain for any $a\in A$:
$\delta(a)$ $=m(\iota\otimes\varepsilon)\Delta\big{(}\delta(a)\big{)}$
$\overset{(\ref{3.4})}{=}m(\iota\otimes\varepsilon)\Big{(}(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}\Big{)}$
$=m(\iota\otimes\varepsilon)\big{(}\delta(a_{(1)})\otimes a_{(2)}+ra_{(1)}\otimes\delta(a_{(2)})\big{)}$
@@
Comparing both sides gives $a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}=0$. Finally,
$\varepsilon\big{(}\delta(a)\big{)}=\varepsilon\Big{(}\delta\big{(}\varepsilon(a_{(1)})a_{(2)}\big{)}\Big{)}=\varepsilon(a_{(1)})\varepsilon\big{(}\delta(a_{(2)})\big{)}=\varepsilon\Big{(}a_{(1)}\varepsilon\big{(}\delta(a_{(2)})\big{)}\Big{)}=0.$
This establishes that $\varepsilon\big{(}\delta(a)\big{)}=0$, confirming the existence of an extension $\varepsilon:A[y;\tau,\delta]\rightarrow k$ that preserves the relation $ya=\tau(a)y+\delta(a)$. For any elements $\sum_{i=1}^{m}a_{i}y^{i},\sum_{j=1}^{n}b_{j}y^{j}\in A[y;\tau,\delta]$, it is straightforward to verify that
$(\varepsilon\otimes\iota)\big{(}\Delta(\sum_{i=1}^{m}a_{i}y^{i})(1\otimes\sum_{j=1}^{n}b_{j}y^{j})\big{)}$ $=(\sum_{i=1}^{m}a_{i}y^{i})(\sum_{j=1}^{n}b_{j}y^{j}),$
$(\iota\otimes\varepsilon)\big{(}(\sum_{j=1}^{n}b_{j}y^{j}\otimes 1)\Delta(\sum_{i=1}^{m}a_{i}y^{i})\big{)}$ $=(\sum_{j=1}^{n}b_{j}y^{j})(\sum_{i=1}^{m}a_{i}y^{i}).$
Thus, we conclude that $\varepsilon:A[y;\tau,\delta]\rightarrow k$ indeed serves as a counit for $A[y;\tau,\delta]$.
@@
Suppose $A[y;\tau,\delta]$ is endowed with the comultiplication and counit as defined above. If an antipode $S$ exists on $A[y;\tau,\delta]$, extended from the underlying algebra $A$ by setting $S(y)=-r^{-1}y$, then $S$ satisfies $(\ref{2.1})$ and $(\ref{2.2})$, while also preserving the relation $ya=\tau(a)y+\delta(a)$. This implies that
@@
$S(a)S(y)=S(y)S\big{(}\tau(a)\big{)}+S\big{(}\delta(a)\big{)}.$ (3.10)
We now proceed to prove equations $(\ref{3.8})$-$(\ref{3.10})$. For $\delta(a)\in A$ it follows from $(\ref{2.1})$ and $(\ref{3.4})$ that
$1\otimes\delta(a)$ $\overset{(\ref{2.1})}{=}(m\otimes\iota)(S\otimes\Delta)\Delta\big{(}\delta(a)\big{)}$
$\overset{(\ref{3.4})}{=}(m\otimes\iota)(S\otimes\Delta)\Big{(}(\delta\otimes\iota)\Delta(a)+(r\otimes 1)\big{(}(\iota\otimes\delta)\Delta(a)\big{)}\Big{)}$
@@
Applying the formula mentioned above, we deduce that for all $a,b\in A$
@@
$\overset{(\ref{3.5})}{=}(m\otimes\iota)(S\otimes\Delta)\Big{(}\tau(a_{(1)})yb\otimes a_{(2)}+\tau(a_{(1)})rb\otimes{a_{(2)}}y+\Delta\big{(}\delta(a)\big{)}(b\otimes 1)\Big{)}$
@@
$\overset{(\ref{3.4})}{=}S\big{(}\tau(a_{(1)})yb\big{)}{a_{(2)(1)}}\otimes a_{(2)(2)}+S\big{(}\tau(a_{(1)})rb\big{)}a_{(2)(1)}y\otimes a_{(2)(2)}$
$\quad+S\big{(}\tau({a_{(1)}})rb\big{)}a_{(2)(1)}r\otimes a_{(2)(2)}y$
$\quad+(m\otimes\iota)\Big{(}S\big{(}\delta(a_{(1)})b\big{)}\otimes a_{(2)(1)}\otimes a_{(2)(2)}+S(ra_{(1)}b)\otimes\delta(a_{(2)(1)})\otimes{a_{(2)(2)}}$
$\quad+S(ra_{(1)}b)\otimes ra_{(2)(1)}\otimes\delta(a_{(2)(2)})\Big{)}$
@@
$\quad+S(b)S(r)S\big{(}\tau(a_{(1)})\big{)}a_{(2)(1)}r\otimes a_{(2)(2)}y+S\big{(}\delta(a_{(1)}b\big{)}a_{(2)(1)}\otimes a_{(2)(2)}$
@@
$\overset{(\ref{3.11})}{=}\big{(}S(b)S(y)+S(b)S(r)y\big{)}\otimes\tau(a)+S(b)\otimes\tau(a)y+(S(b)\otimes 1)(1\otimes\delta(a))$
@@
By comparing both sides of the equation above and using the nondegeneracy of the multiplication in $A$, we conclude that $(\ref{3.8})$ holds. Equation $(\ref{3.9})$ can be verified in a similar manner.
To verify condition $(\ref{3.10})$, we make use of the identities $ya=\tau(a)y+\delta(a)$ and $S(y)=-r^{-1}y$ (where $r^{-1}y\in M(A[y;\tau,\delta])$), which lead to
@@
$=-r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}y-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)}$
for all $a\in A$. Rearranging terms, we obtain
$\bigg{(}r^{-1}\tau\Big{(}S\big{(}\tau(a)\big{)}\Big{)}-S(a)r^{-1}\bigg{)}y=-r^{-1}\delta\Big{(}S\big{(}\tau(a)\big{)}\Big{)}+S\big{(}\delta(a)\big{)}.$
It follows that condition $(\ref{3.10})$ is equivalent to the following pair of identities holding for all $a\in A$: